(-2/11+X=6/11)
I'm pretty sure the answer should be 8/11 but don't know how do the work

Answer by stanbon(75887)

(-2/11)+X=(6/11)
---
Add (2/11) to both sides to get:
X = 8/11
================
cheers,
Stan H.